OpenStudy (shelby1290):

John bought a car for $20 000. He sold it for 30% profit. From the money raised from this slae, he bought another more expensive car, which unfortunately he had to sell for a 30 percent loss. What was his overall profit or loss?

OpenStudy (tkhunny):

You are tempted to say $0, aren't you? That would be incorrect!

OpenStudy (shelby1290):

lol

OpenStudy (shelby1290):

maybe

OpenStudy (shelby1290):

but honestly how do you solve this question

OpenStudy (tkhunny):

Very honest of you. 20000 + 30% profit. What was the sale price of the 1st vehicle?

OpenStudy (shelby1290):

so that would be 20 000 + 30 or 0.3

OpenStudy (shelby1290):

i believe i understand that part

OpenStudy (shelby1290):

so 6000 is the profit and 2600 is the cost for the first car

OpenStudy (tkhunny):

The error was in the notation. Thanks for pointing that out. I fixed it here and deleted the previous post: 20000 + 0.30*20000 = 20000 + 6000 = 26000 OR 20000 * (1.30) = 26000 Make sense?

OpenStudy (shelby1290):

ok i get that

OpenStudy (tkhunny):

Okay, now we buy a new car for 26000. 26000 - 30% or 26000 gives what?

OpenStudy (shelby1290):

25999.7

OpenStudy (tkhunny):

You made the same error that you made on the profit. You cannot just use the percentage. It is 30% OF SOMETHING. In this case, it is 30% OF $26,000. What is 30% of $26,000?

OpenStudy (shelby1290):

so... 30/100x26000 =7800

OpenStudy (tkhunny):

Perfect. So, what was the selling price of the second vehicle?

OpenStudy (shelby1290):

7800

OpenStudy (tkhunny):

No, that's just the loss. It was $26,000 and poor John lost $7800 on the deal. This makes for a total selling price of $26000 - $7800 = ???

OpenStudy (shelby1290):

18200

OpenStudy (tkhunny):

That's it. We are now ready to answer the question. Please reread the problem statement and see what you think about it.

OpenStudy (shelby1290):

Thank you and i believe that is right in my math package it says 1800 but its typed so my teacher mustve made a mistake! anddd another thing, are these chats saved somewhere where i can review them . sorry today is my VERY FIRST time here

OpenStudy (ranga):

1800 is correct. The question asks: What was his overall profit or loss? He made a profit of $6000 on the first car and had a loss of $7800 on the second car. In effect he lost $1800

OpenStudy (shelby1290):

i see, the previous person might have made a mistake and made me do 26000-7800

OpenStudy (shelby1290):

thanks for clearing that up though

OpenStudy (tkhunny):

Alternatively, we can think of it without adding things up. We started with $20,000 worth of car and we ended up with $18,200 worth of cash. $20,000 - $18,200 = $1,800 - Done!

OpenStudy (shelby1290):

Alright so there are two ways but both are correct. I believe i get that

OpenStudy (ranga):

You can go to your profile and click on Question Asked and it will list all the questions you posted. You can choose which one you want to review later on. All these will be saved unless people delete their replies later on.

OpenStudy (shelby1290):

thank you

OpenStudy (ranga):

you are welcome.